OpenStudy (anonymous):

HELPP!!! I need someone to walk me through this problem!! A company placed $1,000,000 in three different accounts. It placed part in short-term notes paying 4.5% per year, twice as much in government bonds paying 5%, and the rest in utility bonds paying 4%. The income after one year was $45,500. How much did the company place in each account?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@xapproachesinfinity @TheSmartOne

jimthompson5910 (jim_thompson5910):

Let x = amount (in dollars) invested in short term notes y = amount (in dollars) invested in government bonds z = amount (in dollars) invested in utility bonds

jimthompson5910 (jim_thompson5910):

"A company placed $1,000,000 in three different accounts" so that means we have this equation x+y+z = 1000000

jimthompson5910 (jim_thompson5910):

we know that "twice as much in government bonds " were invested compared to short term notes, which means we have y = 2x

jimthompson5910 (jim_thompson5910):

so we can go back to the first equation x+y+z = 1000000 and plug in y = 2x x+y+z = 1000000 x+2x+z = 1000000 3x+z = 1000000

OpenStudy (anonymous):

I thought it would be x= 1000000 y= 1000000 z= 1000000 Because the company placed it in three different accounts

jimthompson5910 (jim_thompson5910):

no I think that 1 million dollar figure is the total spread out the 3 accounts

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

If it was like that, then the answer to "How much did the company place in each account?" is simply "1 million dollars each"

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

anyways, "It placed part in short-term notes paying 4.5% per year" so it earns 0.045x dollars from short-term notes ------------------------ "government bonds paying 5%" ---> earns 0.05y dollars from government bonds ------------------------- "utility bonds paying 4%" ---> earns 0.04z dollars with utility bonds

jimthompson5910 (jim_thompson5910):

in total, it earns 0.045x + 0.05y + 0.04z dollars

jimthompson5910 (jim_thompson5910):

"The income after one year was $45,500", so we can say 0.045x + 0.05y + 0.04z = 45500

jimthompson5910 (jim_thompson5910):

Now plug in y = 2x 0.045x + 0.05y + 0.04z = 45500 0.045x + 0.05(2x) + 0.04z = 45500 0.045x + 0.10x + 0.04z = 45500 0.145x + 0.04z = 45500

jimthompson5910 (jim_thompson5910):

after all this, we get this system of equations 3x+z = 1000000 0.145x + 0.04z = 45500

OpenStudy (anonymous):

wait where did 3x come from?

jimthompson5910 (jim_thompson5910):

I plugged y = 2x into x+y+z = 1000000, then combined like terms

OpenStudy (anonymous):

OHH alright

OpenStudy (anonymous):

i can solve it from here thnx!

jimthompson5910 (jim_thompson5910):

tell me what you get so I can check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-0.04*(3x+z = 1000000) 0.145x + 0.04z = 45500 -------------------- -0.12x -0.04 = -40000 0.145x +0.04 = 45500

OpenStudy (anonymous):

-0.12x -0.04z = -40000 0.145x +0.04z = 45500

OpenStudy (anonymous):

0.025x = 5500 x = 220000

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so y = ??

OpenStudy (anonymous):

hold up i got this

OpenStudy (anonymous):

I found z first, though, before getting the y because it was faster. 3x+z = 1000000 3(220000) + z = 1000000 660000 + z = 1000000 z = 340000

jimthompson5910 (jim_thompson5910):

good, and I'm sure you see how to get y from x

OpenStudy (anonymous):

x+y+z = 1000000 220000 + y + 340000 = 1000000 560000 + y = 1000000 y = 440000

jimthompson5910 (jim_thompson5910):

or you can say y = 2x ---> y = 2*(220,000) = 440,000 either way works

OpenStudy (anonymous):

oh yeah that was probably easier :)

jimthompson5910 (jim_thompson5910):

because we let x = amount (in dollars) invested in short term notes y = amount (in dollars) invested in government bonds z = amount (in dollars) invested in utility bonds and because you found that x = 220,000 y = 440,000 z = 340,000 this means $220,000 was invested in short term notes $440,000 was invested in government bonds $340,000 was invested in utility bonds

jimthompson5910 (jim_thompson5910):

you got all 3 correct

OpenStudy (anonymous):

yayy!

OpenStudy (anonymous):

thnx for all the help! i gave u a medal :)

jimthompson5910 (jim_thompson5910):

thx and I'm glad it's clicking now